Technical Report

In the engineering and scientific professions, communication skills are as important as in other fields. Following the development of an innovative design and a cost estimate that predicts large profit, the designer must be able to communicate the findings to other people. There is an old adage that a tree falling in a forest doesn’t make a sound unless there is someone to listen. Similarly, the best technical design in the world might never be implemented unless the designer can communicate the design ideas to the proper people in the right way.

The quality of a report generally provides an image in the reader’s mind that, in large measure, determines the reader’s impression of the quality of work. Of course, an excellent job of report writing cannot disguise a sloppy investigation, but many excellent design studies have not received proper attention and credit because the work was reported in a careless manner.

A formal technical report usually is written at the end of a project. Generally, it is a complete, stand-alone document aimed at persons having widely diverse backgrounds. Therefore, much more detail is required. The outline of a typical formal report consist of

1. Cover page: Should have

• title of the report

• name of author(s)

• address

2. Summary: summarize the work and should also include a short conclusion

3. Table of content

4. List of figures and list of tables include the corresponding page number.

5. Introduction: Back ground of the work (market analysis) to acquaint reader with the problem and purpose of carrying on the work.

6. Design Process: Details of the procedure followed in the design process.

7. Discussion: The discussion could be divided into several subsections such as This section should contain the comprehensive explanation of the results. The discussion could be divided into several subsections such as

• technical analysis (force requirement, speed, etc.)

• details of the build up of the artifact

• equipment used

• assembly and setup procedures

• experimental setups and results

• details of the final results

8. Conclusions: This section states in as concise a form as possible the conclusions that can be drawn from the study.

9. References This section lists all the documents to which the writers referred. The information on each document must be complete and must follow the same formate throughout.

10. Appendices: This includes material deemed to be beyond the scope of the main body of the report. The appendix section may be divided into as many subsections as necessary.

1 Steps in Writing a Report

The five operations involved in the writing of a high-quality report are best remembered with the acronym POWER

1. P: Plan the writing

2. O: Outline the report

3. W: Write

4. E: Edit

5. R: Rewrite

Good writers usually produce good reports. Some of the attributes of good writing are

1. Write as objectively as possible. DO not become emotionally involved or attached to a problem or a solution

2. Be reasonably methodical

3. Record whatever is learned, and keep in mind that whatever work is performed must eventually be documented.

4. Always Strive for clarity in writing, and keep in mind that the written material should be simple and straightforward.

5. Deliver the written material on time. There are also certain qualities associated with good report such as

1. Deliver on due date

2. Effectively answers readers’ questions as they arise

3. Gives a good impression at the first sight

4. Read coherently

5. Contain effective summary and conclusion

6. Is written Clearly and concisely and avoid vague or superfluous phrases 7. Provides pertinent information

2 Illustration Guidelines

visual elements, such as figures, charts, graphs, etc., in technical reports have specific purpose and convey specific information. Visuals are used to explain, illustrate, demonstrate, verify or support written material. Visuals are only valuable if their presentation is effective. The general guidelines for preparing effective visuals (illustrations) are as follows

1. Reference all illustrations in the text 2. Reference the data source

3. Carefully plan the placement of illustrations.

4. Specify all units of measure and the scale used in the drawing

5. Label each illustration with an identifying caption and title. Include figure source, if it was obtained from another document

6. Spell words out rather than using abbreviations

7. When a document has five or more figures, include a list of figures at the beginning of the report.

8. Avoid putting too much on an illustration

3 Mechanics of writing

The following suggestions are presented as a guide to writing and an as in avoiding some of the most common mistakes.

• Paragraph structure: Each paragraph should begin with a topic sentence that provides an overall understanding of the paragraph. Since each paragraph should have a single theme or conclusion, the topic sentence states that theme or conclusion.

• Sentence length: Sentences should be kept as short as possible so that their structure is simple and readable. Long sentences require complex construction, provide an abundance of opportunity for grammatical errors, lake considerable writing time, and slow the reader down. Long sentences are often the result of putting together two independent thoughts that could be stated better in a separate sentences.

• Pronouns: There is no room for any ambiguity between a pronoun and the noun for which it is used. Novice commonly use ”it”, ”this”, ”that”, etc., where it would be better to use one of several nouns. It may be clear to the writer, but it is often ambiguous to the reader. In general, Personal pronouns (I, you, he, she, my, mine, our, us) are not used in technical reports.

• Spelling and punctuation: Errors in these basic elements of writing in the final draft of the report are inexcusable.

• Tense: use the following rules when choosing the tense of the verb when writing

– Past tense: Use to describe work done when you building or designing, or in general to the past events.

– Present tense: Use in reference to items and ideas in the report itself

– Future tense: Use in making prediction from the data or results that will be applicable in the future.

Presentation Style

Unlike the advertising executive, the engineer is ill-equipped to sell his ideas. Second hand information represented by the company officials may not answer all client’s questions. In this write-up and during the semester, students will be given the opportunity to be trained in technical presentations. In this write-up the techniques of oral presentations are briefly discussed.

4 Objective

Every presentation should have an objective. The speaker main objective is to deliver the message (objective) to the audience. The objectives may vary from one talk to another. To identify the real objective one needs to ask the following question: ”If everything goes perfectly, What I intended to achieve?”. One also should realize who are the audience and their educational level. The time in most cases is limited. It is of utmost importance to keep within scheduled time some question/answer session. For this detailed planing is required. Different tools could be used in the presentation such as, slides, models,transparency, audio-visual, and web. Make sure if more than one medium is used, time to shift is accounted for. One easy way to see effectiveness of presentation while in a team is for one to play the role of the speaker and the rest of the team as the audience and possibly play the devils advocate.

4.1 Oral Presentation - Obstacles

To sell your ideas to others, you should first be convinced that your ideas will accomplish the task. Oral presentation requires high degree of creative ability. Recall that humans resist changes though they announce that they welcome it. Humans like familiar methods. Changes requires additional effort which humans in general resist. Some usual quotes one hears:

• We tried before

• It is too radical change

• We have never done that before

• Get back to reality

• We have always done it this way

• I don’t like the idea The speaker need to keep in mind that

• Only 70% of the spoken words is actually received and understood. Complete understanding can come through repetition and redundancy in speech.

• People mostly understand three dimensional objects. Two dimensional projections need to be transmitted with added details.

• Usually people perceive problems from their prospectives.

• Convey ideas so that it is interpreted with the least expenditure of energy.

5 Oral Presentation Dos and Donts

1. Know your audience thoroughly. Deliver an extempore- prepared talk 2. Never read from notes, a sheet or directly from an overhead projector 3. Bring audience up to the speed, in the first few moments 4. Stay within the time allotted

5. Include relevant humorous stories, anecdotes or joke ( Only if you are good at it)

6. Avoid using specialized technical jargon. Explain the terms if you feel the audience may not know.

7. Understand your message clearly. The whole goal is to clearly communicate the message

8. Practice-practice-practice, may like to memorize introduction and concluding remarks

9. The dry run is a dress rehearsal. Use it to iron out problems in delivery,organization and timing.

10. Avoid mannerisms: Speak confidently but not aggressively

11. Maintain eye-to-eye contact with some audience and keep shifting that contact through the talk.

12. Never talk to the board or to empty space

13. Present the material in a clever fashion, but not in a cheap and sensational fashion. Be genuinely sincere and professional

14. Logical presentation is much more critical in oral than in written presentation

6 Oral Presentation Techniques

1. Visual aids (sketches, graphs, drawings, photos, models, slides, transparency, web) often convey information more efficiently and effectively. Visual aids permits dual sense of hearing and seeing, apart from helping the speaker.

2. Limit slides (Viewgraphes) to not more than one per minute

3. each slide should contain one idea 4. First Slide should show the title of your talk and names.

5. Second slide should give a brief outline of the talk

6. Last slide should summarize the message you just delivered

7. If you need to show a slide more than once use a second copy

8. Avoid leaving a slide on the screen if you have finished discussion on that topic.

9. Never read directly from the slide. Spoken words should complement the slides. Prepare notes for each slide and use it during practice.

10. Use graphs to explain variations. Clearly label the axis, data, and title. acknowledge the source.

11. Every graph should have a message ( idea). Color should enhance the communication not distract.

12. Audience respond to well organized information. That means, (a) Efficient presentation

(b) All assumptions clearly stated and justified

(c) Source of information and facts are clearly outlined

13. Begin with the presentation of the problem and conclusion/recommendation (primarily goal)

14. Finish ahead and be prepared for the question answer session.

7 Question/Answer session

The question answer session is very important. It shows the enthusiasm in the audience and usually shows interest and attention. In the Q/A session you should

1. Allow the questioner to complete the question, before answering

2. Avoid being argumentative 3. Do not let the questioner feel that the question is stupid 4. Adjourn the meeting, if the questions slack off

5. Thank the audience, one final time after the Q/A session